What is the acceleration of a proton moving with a speed of 8.0 m/s at right angles to a magnetic field of 1.5 T ?
Answer #1

Using force balance

Fnet = Fm

m*a = q*VxB

angle between V and B = 90 deg

VxB = V*B*sin 90 deg = V*B

m*a = q*V*B

a = q*V*B/m

Using given values:

a = 1.6*10^-19*8*1.5/(1.67*10^-27)

a = 1.15*10^9 m/sec^2

Please Upvote.